# Style over Substance: Failure Modes of LLM Judges in Alignment Benchmarking

### v1

@@ Abstract
The release of ChatGPT in November 2022 sparked an explosion of interest in post-training and an avalanche of new preference optimization (PO) methods. These methods claim superior alignment by virtue of better correspondence with human pairwise preferences, often measured by LLM judges. In this work, we attempt to answer the following question – do LLM-judge preferences translate to progress on other, more concrete metrics for alignment, and if not, why not? We define a concrete metric for alignment, and introduce SOS-Bench, the largest standardized, reproducible LLM meta-benchmark to date. We find that (1) LLM-judgments do not correlate with concrete measures of safety, world knowledge, and instruction following; (2) LLM judges have powerful implicit biases, prioritizing style over factuality and safety; and (3) the supervised fine-tuning (SFT) stage of post-training, and not the PO stage, has the greatest impact on alignment, with data scaling and prompt diversity as the driving factors. Our codebase and complete results can be found at https://anonymous.4open.science/r/mismo-bench-587D/readme.md.
@@ 1 Introduction
The release of ChatGPT in November 2022 sparked an explosion of interest in post-training and an avalanche of new preference optimization (PO) methods and data curation strategies for supervised fine tuning (SFT). Many of these recent works evaluate primarily or exclusively using LLM-judge benchmarks such as MT-Bench, Alpaca Eval, and Arena-Hard-Auto (Dubois et al., 2024; Li et al., 2024c; Zheng et al., 2023). Such benchmarks employ LLM judges that are intended to simultaneously automate evaluation while also aligning with human preference.
@@
  • We find that LLM judges have powerful implicit biases. Specifically, they prioritize stylistic preferences over other important considerations in alignment, like factuality and safety.
@@
Figure 1: The LLM-judge pipeline introduces new potential confounds in evaluation, compared to standard benchmarks. We diagram the LLM-judge pipeline and observe that it is more complex than that of standard benchmarks, raising the potential for new confounding variables. Rather than relying on an objective ground truth, preference benchmarks substitute the preferences of a judge (in this case, also an LLM). This introduces new potential confounds: (1) the choice of judge, (2) the instructions to the judge, and (3) implicit biases which can affect a judge’s stated preferences independent of any instructions.
@@ 2 Preliminaries
3 LLM-Judge Benchmarks
Post-training methods papers are increasingly focused on LLM-judge benchmarks. We motivate our work by establishing how widespread the practice of reporting solely or primarily on LLM-judge benchmarks has become since 2022. We select, using search terms that are in common use in the field, a representative sample of 19 papers from the literature, all but one of which were published after the debut of ChatGPT (8 from 2023, 10 from 2024). In addition, we survey the core claims of these papers, as referenced in their abstracts and introductions. Details about the methods assessed can be found in Table 7.
We find that 80% of methods benchmark exclusively on pairwise preferences using either LLM or human judges, up from 50% in 2023, with 70% only using LLM judges, up from 37.5% in 2023. This abrupt surge in use of these benchmark approaches invites closer consideration of their properties and design.
@@ 3 LLM-Judge Benchmarks
LLM-judge scores are inconsistent with static benchmarks, but consistent with each other. LLM-judge benchmarks are designed to closely mimic pairwise human preferences. However, it is unclear what these benchmarks actually measure. To address this question, in Table 1 we compare the ranking of eight top-performing LLMs on LiveBench, HELM, Arena-Hard-Auto, and ChatBot Arena (White et al., 2024; Liang et al., 2023). We find that LLM-judges indeed have preferences that closely (although imperfectly) correlate with those of humans, but their correlation with static benchmarks is weak. In order to understand why this might be the case, we introduce a framework for interpreting and comparing them to one another.
@@
Towards a universal framework for LLM judges. We observe that all pipelines which employ LLMs as surrogates for human pairwise preferences necessarily make use of certain components. In order to facilitate analysis of these systems, we diagram their common architectural flow in Figure 1.
@@
From this diagram, it quickly becomes clear that there are several key structural distinctions to be made between LLM-judge benchmarks and traditional NLP benchmarks, many of which introduce new confounds into the evaluation process:
  • Traditional benchmark metrics are model-free; LLM-judge benchmarks require a judge. The choice of judge is a significant confound, but it is not yet standard practice to ablate this choice, or ensemble across judges.
@@
  • Traditional benchmark scores are reference-based; LLM-judge benchmarks are comparison-based. The choice of baseline response to which assistants are compared represents yet another potential confound, as pairwise preferences over texts do not obey the transitive property.
@@
  • Traditional benchmarks contain many questions on a narrow subject; LLM-judge benchmarks contain a small number of questions on a wide range of subjects. Because of their high unit costs, LLM-judge benchmarks are smaller than standard NLP benchmarks like GLUE and MMLU; Arena-Hard has a test set of 500 questions. The authors have justified this choice by demonstrating that their benchmarks nevertheless correlate strongly with preference reports on ChatBot Arena. But this does not guarantee that a preference score on a broad range of topics and people will correlate well with individual use cases, as capabilities are vector-valued, not scalar-valued. To the extent that strong correlations are achievable, they would only be so if the judgment criteria were consistent across all tasks and all judges, which would tend to favor stylistic factors.
@@
  • Traditional benchmarks specify a metric, but LLM-judge benchmarks must specify both a metric and judging criteria. This introduces a new potential confound not present in traditional benchmarks; the instructions to the judge may be underspecified, unclear, or may simply not reflect best practices in model alignment with human preference.
@@
Table 1: Pairwise preference benchmarks do not track established benchmarks. We report Pearson’s R as a measure of the strength of correlation between pairs of traditional benchmarks and LLM-judge benchmarks. LiveBench and HELM are strongly correlated, as are Arena-Hard and ChatBot Arena. All other pairs show comparatively weaker correlation.
@@ 5 SOS-Bench
SOS-Bench combines 19 existing world knowledge, instruction following, and safety benchmarks for a holistic view of model performance. The benchmark name is derived from the phrase More is More in alignment, which is a succinct summary of our findings in Section 6. For the complete list of benchmarks we use, please refer to Table 8. All of the questions in our benchmark contain ground truth answers, and aggregates are reported using the average of normalized accuracies (by the size of the test set), with 95% confidence intervals.
Comparing SOS-Bench to existing evaluations. All in all, we test models on 152,380 data points; to the best of our knowledge, this is almost 7x larger than the largest previous open source LLM benchmark which end users can run themselves, HuggingFace’s OpenLLM Leaderboard 2. While individual model releases and technical reports also release meta-benchmark results, they suffer from two failings; they are generally not reproducible, and the aggregations of results, which are different for each model release, are vulnerable to cherry picking. By combining scale and standardization, we hope to create a reliable, concrete alignment measure.
@@
Measuring alignment. We subdivide our benchmark into three factors (world knowledge, instruction following, and safety) and report the results for each. For the sake of comparison, we also report results from Arena-Hard-Auto. For results on individual benchmarks, we refer the reader to https://anonymous.4open.science/r/mismo-bench-587D/readme.md.
@@ 8 Related Work
A series of recent works have offered various critiques on the use of LLM-judges (Chen et al., 2024; Bavaresco et al., 2024; Wei et al., 2024). Several known confounds for pairwise benchmarks are well-established in the literature, including a bias by both humans and LLMs in favor of longer responses and a preference of LLMs for the style of their own output, and efforts have been made to control for length as a confound (Panickssery et al., 2024; Park et al., 2024; Dubois et al., 2024). There also exists considerable prior literature critiquing human judgment bias and proposing mechanisms for collective decision making, including social choice theory and prospect theory (Ge et al., 2024; Ethayarajh et al., 2024). Finally, there exists a literature on the limitations of RLHF and pairwise preferences, including findings that universal AI alignment using RLHF is impossible (Singhal et al., 2024; Casper et al., 2023; Lambert & Calandra, 2024; Mishra, 2023; Lambert et al., 2023). Extending prior work, our research focuses on novel semantic biases, and also introduces the concept of an implicit weighting of factors on the part of LLM judges. Finally, our work draws on all of the aforementioned traditions to produce a meta-analysis on the progress, or lack thereof, of methods for LLM post-training.
@@ 9 Impact / Limitations
Our work advocates for the development of explicit inductive bias in LLM-judges. If implemented naively, this could lead to Goodhart effects. It is therefore essential that special research emphasis be placed on the development and standardization of judging templates and metrics, and that these factors be continuously optimized to prevent overfitting (Manheim & Garrabrant, 2019).
@@ Appendix C Method Comparison
Table 7: Recent work in post-training has relied heavily on LLM judgments. Post-training stage (PT) value 1 refers to SFT datasets and methods, 2 refers to preference optimization methods, 3 refers to other methods. For the Only LLM-Judge and Only PO columns, 1 is positive and 0 is negative.
@@
Zephyr from Tunstall et al. (2023) 2023 2 1 1
WizardLM from Xu et al. (2024a) 2023 1 1 1

### v2

# Style Outweighs Substance: Failure Modes of LLM Judges in Alignment Benchmarking

@@ Abstract
The release of ChatGPT in November 2022 sparked an explosion of interest in post-training and an avalanche of new preference optimization (PO) methods. These methods claim superior alignment by virtue of better correspondence with human pairwise preferences, often measured by LLM-judges. In this work, we attempt to answer the following question – do LLM-judge preferences translate to progress on other, more concrete metrics for alignment, and if not, why not? We define a concrete metric for alignment, and introduce SOS-Bench (Substance Outweighs Style Benchmark), the largest standardized, reproducible LLM meta-benchmark to date. We find that (1) LLM-judge preferences do not correlate with concrete measures of safety, world knowledge, and instruction following; (2) LLM-judges have powerful implicit biases, prioritizing style over factuality and safety; and (3) the supervised fine-tuning (SFT) stage of post-training, and not the PO stage, has the greatest impact on alignment, with data scaling and prompt diversity as the driving factors. Our codebase and complete results can be found at https://github.com/penfever/sos-bench/.
@@ 1 Introduction
The release of ChatGPT in November 2022 sparked an explosion of interest in post-training and an avalanche of new preference optimization (PO) methods and data curation strategies for supervised fine tuning (SFT). Many of these recent works evaluate primarily or exclusively using LLM-judge preference benchmarks such as MT-Bench, Alpaca Eval, and Arena-Hard-Auto (Dubois et al., 2024; Li et al., 2024c; Zheng et al., 2023). Such benchmarks employ LLM-judges that are intended to automate evaluation and align with human preferences.
@@
  • We find that LLM-judges for alignment benchmarking have powerful implicit biases. Specifically, they prioritize stylistic preferences over other important considerations, like factuality and safety.
@@
Figure 1: The LLM-judge pipeline introduces new potential confounds in evaluation, compared to standard benchmarks. We diagram the LLM-judge pipeline for alignment benchmarking and observe that it is more complex than that of most standard benchmarks; (a) it replaces an explainable, deterministic metric with an opaque LLM-judge. (b) it does not attempt to establish any verifiable ground truth. (c) it contains a relatively small number of questions covering an very wide range of topics, resulting in limited coverage of any particular knowledge domain. (d) it introduces novel confounds in the form of the judging template (explicit bias) and the judge’s unstated internal preferences (implicit bias).
@@ 2 Preliminaries
3 LLM-Judge Preference Benchmarks
LLM-judge preference benchmarks (hereafter referred to as LLM-judge benchmarks), such as Arena-Hard-Auto, Alpaca Eval, and MT Bench, have become quite popular since their introduction shortly after the initial release of Llama. Today, many papers report solely or primarily on these benchmarks. This abrupt surge in their use invites closer consideration of their properties and design.
@@ 3 LLM-Judge Preference Benchmarks
Pairwise preference benchmark scores are inconsistent with static benchmarks, but consistent with each other. In Table 1 we compare the ranking of eight top-performing LLMs on LiveBench, HELM, Arena-Hard-Auto, and ChatBot Arena (White et al., 2024; Liang et al., 2023). We find that LLM-judges of alignment indeed have preferences that closely (albeit imperfectly) correlate with those of humans, but that their correlation with static benchmarks is weak. Similar effects can be observed on the leaderboard of BenchBench, a recent paper which aims to standardize benchmark agreement testing.  (Perlitz et al., 2024). When we aggregate using standard benchmarks (HELM, HuggingFace OpenLLM Leaderboard 2, LiveBench and OpenCompass), the highest overall BenchBench score is 2.2, and the highest pairwise preference score is 0.69. Conversely, if we instead aggregate using preference benchmarks (Alpaca Eval, MT-Bench, Arena-Hard, Chatbot Arena) the highest overall score is 1.8, and the highest standard benchmark score is 1.4. Such measures, however, cannot tell us which benchmark regime we ought to trust more, or why (Perlitz et al., 2024). In order to understand this point better, we introduce a framework for interpreting and comparing them to one another.
@@
Towards a universal framework for LLM-judge benchmarks. We observe that all pipelines which employ LLMs as surrogates for human pairwise preferences necessarily make use of certain components. In order to facilitate analysis of these systems, we diagram their common architectural flow in Figure 1.
@@
From this, it quickly becomes clear that there are several key structural distinctions to be made between LLM-judge benchmarks and standard NLP benchmarks:
  • Most standard benchmark metrics are model-free; all LLM-judge benchmarks require a judge model. The most commonly used standard benchmarking metrics, such as BLEU, are model-free. While both the choice of metric and the choice of judge can be a potential confound, the opacity of LLMs makes their behavior much more challenging to interpret, compared to deterministic model-free metrics. Despite this, it is not yet standard practice to ablate this choice, or ensemble across judges, most likely for reasons of cost.
@@
  • Standard benchmark scores are reference-based; LLM-judge benchmarks are comparison-based. The choice of baseline response to which assistants are compared represents another potential confound, as pairwise preferences over texts do not obey the transitive property. [^2] It is worth noting that reference-based labels also can fall short of the idealized ground truth, in particular on open-ended generative tasks such as summarization.
@@
  • Standard benchmarks contain many questions on a narrow subject; LLM-judge benchmarks contain a small number of questions on a wide range of subjects. Because of their high unit costs, LLM-judge benchmarks are smaller than standard NLP benchmarks like GLUE and MMLU; Arena-Hard has a test set of 500 questions. The authors have justified this choice by demonstrating that their benchmarks nevertheless correlate strongly with preference reports on ChatBot Arena. But this does not guarantee that a preference score on a broad range of topics and people will correlate well with individual use cases, as capabilities are vector-valued, not scalar-valued. To the extent that strong correlations are achievable, they would only be so if the judgment criteria were consistent across all tasks and all judges, which would tend to favor stylistic factors.
@@
  • Standard benchmarks specify a metric, but LLM-judge benchmarks must specify both a metric and judging criteria. This introduces a new potential confound not present in standard benchmarks; the instructions to the judge may be underspecified, unclear, or may simply not reflect best practices in model alignment with human preference.
@@
Table 1: Pairwise preference benchmarks do not track established benchmarks. We report Pearson’s R as a measure of the strength of correlation between pairs of traditional benchmarks and pairwise preference benchmarks. LiveBench and HELM are strongly correlated, as are Arena-Hard and ChatBot Arena. All other pairs show comparatively weaker correlation.
@@ 5 SOS-Bench
SOS-Bench (Substance Over Style Benchmark) combines 19 existing world knowledge, instruction following, and safety benchmarks for a holistic view of model performance. For the complete list of benchmarks we use, please refer to Table 8. All of the questions in our benchmark contain ground truth answers, and aggregates are reported using the average of normalized accuracies (by the size of the test set), with 95% confidence intervals.
Comparing SOS-Bench to existing evaluations. All in all, we test models on 152,380 data points; to the best of our knowledge, this is almost 7x larger than the largest previous open source LLM benchmark which end users can run themselves, HuggingFace’s OpenLLM Leaderboard 2. While individual model releases and technical reports also release meta-benchmark results, they suffer from two failings; they are not always reproducible, and the aggregations of results, which are different for each model release, are vulnerable to cherry picking. By combining scale and standardization, we hope to create a reliable, concrete alignment measure.
@@
Measuring alignment. We subdivide our benchmark into three factors (world knowledge, instruction following, and safety) and report the results for each. For the sake of comparison, we also report results from Arena-Hard-Auto. For results on individual benchmarks, we refer the reader to https://github.com/penfever/sos-bench/.
@@ 8 Related Work
A series of recent works have offered various critiques on the use of LLM-judges (Chen et al., 2024; Bavaresco et al., 2024; Wei et al., 2024). Several known confounds for pairwise benchmarks are well-established in the literature, in particular a bias by both humans and LLMs in favor of longer responses and a preference of LLMs for the style of their own output; efforts have been made to control for length as a confound (Panickssery et al., 2024; Park et al., 2024; Dubois et al., 2024; Wu & Aji, 2023). There also exists considerable prior literature critiquing human judgment bias and proposing mechanisms for collective decision making, including social choice theory and prospect theory (Ge et al., 2024; Ethayarajh et al., 2024). Finally, there exists a literature on the limitations of RLHF and pairwise preferences, including findings that universal AI alignment using RLHF is impossible (Singhal et al., 2024; Casper et al., 2023; Lambert & Calandra, 2024; Mishra, 2023; Lambert et al., 2023). Extending prior work, our research focuses on novel semantic biases, and also introduces the concept of an implicit weighting of factors on the part of LLM judges. Finally, our work draws on all of the aforementioned traditions to produce a meta-analysis on the progress, or lack thereof, of methods for LLM post-training.
@@ 9 Impact / Limitations
Our work advocates for the development of explicit inductive bias in LLM-judges. If implemented naively, this could lead to Goodhart effects. It is therefore essential that special research emphasis be placed on the development and standardization of judging templates and metrics, and that these factors be continuously optimized to combat overfitting (Manheim & Garrabrant, 2019).
@@ Appendix B Compute and Resources
The experiments for this work were conducted using NYU’s Greene cluster and on Google Cloud, with additional financial support provided by Arthur.AI.
@@ Appendix C Method Comparison
Table 7: Recent work in post-training has relied heavily on LLM judgments. Post-training stage (PT) value 1 refers to SFT datasets and methods, 2 refers to preference optimization methods, 3 refers to other methods. For the Only LLM-Judge and Only PO columns, 1 is positive and 0 is negative. This table covers the main paper only, and does not include appendix experiments.
@@
Zephyr from Tunstall et al. (2023) 2023 2 0 0
WizardLM from Xu et al. (2024a) 2023 1 0 1